I need help this is the last question on my homework and i have been siting here for 30 mins trying to figure it out. he wants us to find x1 and x2 i factored it and got 4, -7 and its wrong someone help me.
Need to find x1 and x2
Solve the equation x^2 - 3 x - 28=0 by factoring.

Answer by jim_thompson5910(35256)

If you're looking for the solutions (roots or zeros) then...

x^2 - 3x - 28 = 0

(x - 7)(x + 4) = 0

x - 7 = 0 or x + 4 = 0

x = 7 or x = -4

So the solutions are x = 7 or x = -4